# Mixture-of-RAG: Integrating Text and Tables with Large Language Models

### v2

@@ 3.1.1 General RCL Table Summary
$\left\{\begin{aligned} H_{t}&=\{h_{t}(1),h_{t}(2),\dots,h_{t}(n)\mid n\text{ is \# columns}\}\\ H_{l}&=\{h_{l}(1),h_{l}(2),\dots,h_{l}(m)\mid m\text{ is \# rows}\}\end{aligned}\right$. (5)
@@
$\left\{\begin{aligned} S_{\text{row}}\{h_{l}(i)\}&=\mathcal{G}(h_{l}(i),\bigcup_{j=1}^{n}h_{t}(j))\\ S_{\text{col}}\{h_{t}(j)\}&=\mathcal{G}(h_{t}(j),\bigcup_{i=1}^{m}h_{l}(i))\end{aligned}\right$. (6)
, where $\mathcal{G}$ denotes the operation that generates the general RCL summary for both rows and columns. This operation condenses the information in each row and column, simplifying the table’s hierarchical structure for retrieval.
3.1.2 H-RCL Table Summary
@@
$\left\{\begin{aligned} S_{\text{row}}(h_{l}^{r}(i))&=\mathcal{G}(P_{l}^{r}(i),\bigcup_{j=1}^{n}P_{t}^{R_{t}}(j))\\ S_{\text{col}}(h_{t}^{r}(j))&=\mathcal{G}(P_{t}^{r}(j),\bigcup_{i=1}^{m}P_{l}^{R_{l}}(i))\end{aligned}\right$. (7)
@@
$S_{\text{table}}=\bigcup_{i,r}S_{\text{row}}(h_{l}^{r}(i))\cup\bigcup_{j,r}S_{\text{col}}(h_{t}^{r}(j))$ (8)
@@
3.1.3 Passage Processing
In addition to processing tables, we also address the handling of text passages within the heterogeneous document. In this module, we focus on processing the passage at the sentence level, dividing the passage into individual sentences, and constructing the corpus. We segment the passage at the sentence level to enable more effective handling of textual data, ensuring that relevant information is preserved and structured for retrieval and reasoning.
Figure 3: The Path and Hierarchical Levels in the Table. The H-RCL Table Summary Relies on the Structure and Paths of Hierarchical Table
3.2 Cross-Modal Retrieval Module
@@
The ensemble approach combines the strengths of BM25 and embedding-based retrieval to create a diverse candidate document set. BM25 ensures accurate keyword-based matches, while embedding-based retrieval adds semantic depth and flexibility. By merging, the ensemble retrieval stage provides high recall and relevance, laying a strong foundation for the subsequent LLM-based retrieval stage.
@@
$S=\text{SelectTop}(Q,D,\Theta)$ (9)
@@
Subsequently, for each document $D_{i}$ in the top-$k$ documents, only the information belonging to this collection is retained, and the filtered information is recombined into a "new document" $D^{\prime}_{i}$ to replace the original document:
$D^{\prime}_{i}=\text{Recombine}(S,D_{i})$ (10)
@@
3.3 Multi-Step Reasoning Module
@@
We execute QA tasks as an instance based on the most relevant document $D^{*}$ provided by the Cross-Modal Retrieval Module. Although LLMs demonstrate strong performance in natural language understanding, performing QA on heterogeneous documents remains a challenging task. To more accurately extract relevant information from heterogeneous documents and leverage the information to answer the questions, we introduce a novel prompting strategy for QA tasks with MixRAG called RECAP (Restate, Extract, Compute, Answer, Present), inspired by prompt strategies such as CoT [25] and EEDP [19] that involve step-by-step analytical reasoning, as well as PoT [6] which generates executable programs to solve complex mathematical operations.
@@
Formally, as defined in Section 2.1, the Multi-Step Reasoning Module is represented as $A=\text{Inference}(Q,D^{*})$, where $A$ is the final answer produced by the RECAP strategy.
The specific workflow of RECAP is as follows: (1) Restate the Question: Clearly restate the requirements of the question and analyze the question type; (2) Extract Relevant Data: Extract and list all the information from the heterogeneous document that is directly relevant to the question; (3) Perform Logical Reasoning: The LLM is required to perform step-by-step reasoning to obtain the final answer, and the LLM should also provide the complete calculation formula used to derive the final result (if the question is not calculation-based, the result is provided directly); (4) Provide Final Answer: For the answer derived by the LLM through reasoning and the other answer obtained by the external calculator based on the provided formula, we employ a rule-based approach to select the more appropriate answer from the two, depending on the question type, as the final answer. Unlike methods like Self-Consistency [23], which improve accuracy by generating multiple reasoning paths, RECAP adopts a single-turn dialogue strategy, with its logical reasoning and formula provision originating from the same reasoning path, offering advantages in token efficiency for the LLM. The comparison between RECAP and various QA inference baselines is given in Appendix  G.
4 Experiment
4.1 Experimental Setup
To provide a comprehensive evaluation of MixRAG, we conduct experiments comparing its performance with relevant baseline methods. Detailed descriptions of these baselines are provided in Appendix C. We adopt HiT@K and Exact Match (EM) as evaluation metrics for retrieval and QA, respectively. Further details are available in Appendix D. We employ GPT-4o throughout our pipeline, including heterogeneous document representation, LLM-based retrieval, and RECAP execution. For embedding-based retrieval, we use OpenAI’s text-embedding-3-large model as the encoder. According to Appendix F, we set the parameters of MixRAG to $n=40$ and $m=60$ for ensemble retrieval. For the QA instance RECAP, the number of examples for all prompt strategies is uniformly set to 1. We evaluate the RECAP instance using GPT-4o, GPT-4o mini, and three open-source large language models: Qwen2.5-32B-Instruct, Mistral-Nemo-Instruct-2407, and Llama-3.1-8B-Instruct. Detailed evaluation results are presented in Appendix G.
4.2 Comparison with Baselines
In this section, we present a comprehensive analysis of MixRAG compared with several strong retrieval-augmented baselines, across the DocRAGLib dataset to evaluate both retrieval performance (HiT@K) and downstream QA accuracy (EM).
@@ 4.2 Comparison with Baselines
Table 1: Comparison of MixRAG Performance with Various Baselines
HiT@K EM*
Method K=1 K=3 K=5 K=10
Standard RAG 0.0159 0.0339 0.0538 0.1255 0.0092
LangChain 0.2390 0.4104 0.4821 0.5219 0.1301
Self-RAG 0.2829 0.4502 0.5000 0.5598 0.2141
Table Retrieval 0.3705 0.5299 0.5996 0.6494 0.2541
MixRAG 0.5410 0.7244 0.7603 0.8689 0.3228
@@
Retrieval Performance Comparison with Baselines. Table 1 presents the retrieval performance of MixRAG compared to several baselines. Across all HiT@K metrics (K=1, 3, 5, 10), MixRAG consistently achieves the best results, reaching a HiT@1 score of 0.541, significantly outperforming the strongest baseline by 46%. The fine-grained row-and-column-level summarization, combined with the two-stage retrieval strategy, achieves substantial improvements in retrieval accuracy for complex heterogeneous documents.
Document-QA Accuracy Comparison with Baselines. Using RECAP as the QA instance module for each retrieval method, we evaluate document-level QA accuracy based on EM scores. According to the EM results in Table 1, MixRAG also achieves the highest QA accuracy, with an EM score of 0.3228, which corresponds to a 27.1% improvement over the second-best baseline, Table Retrieval. These results highlight the effectiveness of our retrieval and reasoning framework in improving answer accuracy for heterogeneous documents retrieval in downstream QA tasks.
4.3 Ablation Study
@@
Impact of Heterogeneous Document Representation Strategy in MixRAG. To evaluate the impact of different table summarization methods in heterogeneous document representation module, we compare three variants: Table Level Summary, General RCL Summary, and H-RCL Summary. As shown in Table 2, the H-RCL Table Summary strategy achieves the best performance among all heterogeneous document representation strategies. The General RCL Summary ranks second, showing clear gains over the traditional Table Level Summary by providing more explicit representations of rows and columns. H-RCL improves HiT@1 by 47% and EM by 28% compared to the Table Level Summary, demonstrating the importance of capturing multi-level table structures. The General RCL strategy provides a clearer representation of rows and columns, reducing ambiguity and improving retrieval accuracy, which in turn enhances QA performance by enabling more precise document matching. Building on this, the H-RCL strategy further captures multi-level table structures and incorporates hierarchical path information, significantly boosting both retrieval accuracy and the model’s ability to perform complex reasoning over structured data. Overall, these results demonstrate the effectiveness of the H-RCL design in improving both retrieval and reasoning capabilities for heterogeneous documents.
Table 2: Performance Evaluation of MixRAG with Different Table Summary Level
heterogeneous document representation HiT@K EM*
Table Summary Level K=1 K=3 K=5 K=10
Table Level Summary 0.3627 0.4805 0.5963 0.7428 0.2383
General RCL Summary 0.3984 0.5502 0.6803 0.7981 0.2625
H-RCL Summary 0.5410 0.7244 0.7603 0.8689 0.3228
*Note: EM scores are based on HiT@1 retrieval using RECAP.
Impact of Retrieval and QA Module. To investigate the effectiveness of each key component in MixRAG, we conduct the ablation study using DocRAGLib dataset. Table 4.3 shows the results of the Retrieval and QA module ablation study. Replacing Ensemble Retrieval with either Embedding or BM25 significantly reduces HiT@1, with Embedding achieving better results due to its stronger semantic alignment capabilities. Further removing LLM-based retrieval leads to an additional decline of 0.19 and 0.20 in HiT@1 for Embedding and BM25, respectively, confirming the advantage of the two-stage retrieval framework over single-stage methods. Moreover, disabling external calculation in RECAP results in a noticeable drop in EM, underscoring the importance of precise formula-based reasoning for improving answer accuracy. Overall, these results validate the effectiveness of each component within MixRAG in enhancing both retrieval precision and answer accuracy.
Table 3: Ablation Study Result of MixRAG
HiT@1 EM
Ours (MixRAG) 0.5410 0.3238
@@
only Embedding 0.4734 0.2777
only BM25 0.3545 0.2100
@@
only Embedding 0.2787 0.2131
only BM25 0.1557 0.1311
- w/o Calc of RECAP 0.5410 0.2838

### v3

@@ 4.1.1. General RCL Table Summary
(5) $\left\{\begin{aligned} H_{t}&=\{h_{t}(1),h_{t}(2),\dots,h_{t}(n)\mid n\text{ is \# columns}\}\\ H_{l}&=\{h_{l}(1),h_{l}(2),\dots,h_{l}(m)\mid m\text{ is \# rows}\}\end{aligned}\right$.
@@
(6) $\small\left\{\begin{aligned} S_{\text{row}}\{h_{l}(i)\}&=\mathcal{G}(h_{l}(i),\bigcup_{j=1}^{n}h_{t}(j))\\ S_{\text{col}}\{h_{t}(j)\}&=\mathcal{G}(h_{t}(j),\bigcup_{i=1}^{m}h_{l}(i))\end{aligned}\right$.
where $\mathcal{G}$ denotes the operation that generates the general RCL summary for both rows and columns. This operation condenses the information in each row and column, simplifying the table’s hierarchical structure for retrieval.
4.1.2. H-RCL Table Summary
@@
(7) $\small\left\{\begin{aligned} S_{\text{row}}(h_{l}^{r}(i))&=\mathcal{G}(P_{l}^{r}(i),\bigcup_{j=1}^{n}P_{t}^{R_{t}}(j))\\ S_{\text{col}}(h_{t}^{r}(j))&=\mathcal{G}(P_{t}^{r}(j),\bigcup_{i=1}^{m}P_{l}^{R_{l}}(i))\end{aligned}\right$.
@@
(8) $S_{\text{table}}=\bigcup_{i,r}S_{\text{row}}(h_{l}^{r}(i))\cup\bigcup_{j,r}S_{\text{col}}(h_{t}^{r}(j))$
@@
4.1.3. Passage Processing
In addition to processing tables, we also address the handling of text passages within the heterogeneous document. In this module, we focus on processing the passage at the sentence level. To ensure that each sentence carries complete semantic meaning, we first apply coreference resolution using spaCy, replacing referential expressions with their explicit noun phrase antecedents. Then the document text is segmented into individual sentences using spaCy’s sentence tokenizer. To reduce noise, we filter out overly short sentences (less than 5 words), which often lack meaningful content. The resulting collection of meaningful sentences forms the textual portion of document corpus, supporting retrieval and downstream reasoning tasks.
Figure 4. The Path and Hierarchical Levels in the Table
4.2. Cross-Modal Retrieval Module
@@
The ensemble approach combines the strengths of BM25 and embedding-based retrieval to create a diverse set of candidate documents. BM25 ensures accurate keyword-based matches, while embedding-based retrieval adds semantic depth and flexibility. By merging, the ensemble retrieval stage provides high recall and relevance, laying a strong foundation for the subsequent LLM-based retrieval stage.
@@
(9) $S=\text{SelectTop}(Q,D,\theta)$
@@
Subsequently, for each document $D_{i}$ in the top-$k$ documents, only the information belonging to this collection is retained, and the filtered information is recombined into a “new document” $D^{\prime}_{i}$ to replace the original document:
(10) $D^{\prime}_{i}=\text{Recombine}(S,D_{i})$
@@
4.3. Multi-Step Reasoning Module
Table 1. Performance Comparison with Various Baselines
Method HiT@K(%) EM*(%)
   (- LLM) K=1 K=3 K=5 K=10
[…]
Langchain
- GPT-4o-mini 20.70 35.55 46.72 49.39 7.89
- GPT-4o 23.90 41.04 48.21 52.19 13.01
- Mistral-Nemo 17.93 31.15 43.55 48.05 7.07
Self-RAG
- GPT-4o-mini 26.33 41.29 48.67 56.45 20.70
- GPT-4o 28.29 45.02 50.00 55.98 21.41
- Mistral-Nemo 19.88 37.19 40.68 54.30 15.58
RAG-Critic
- GPT-4o-mini 25.41 42.93 47.75 59.63 22.23
- GPT-4o 29.00 46.11 50.51 60.55 25.31
- Mistral-Nemo 21.31 37.40 40.88 55.74 19.36
MixRAG
- GPT-4o-mini 48.46 69.36 73.87 84.84 29.51
- GPT-4o 54.10 72.44 76.03 86.89 32.28
- Mistral-Nemo 44.47 67.01 70.08 82.38 27.25
*: EM scores are based on HiT@1 retrieval used to augment answer generation. Standard RAG and Table Retrieval do not use LLMs during retrieval. $\spadesuit$: These EM score reported in order: GPT-4o-mini, GPT-4o, Mistral-Nemo.
@@ 4.3. Multi-Step Reasoning Module
We execute QA tasks as an instance based on the most relevant document $D^{*}$ provided by the Cross-Modal Retrieval Module. Although LLMs demonstrate strong performance in natural language understanding, performing QA on heterogeneous documents remains a challenging task. To more accurately extract relevant information from heterogeneous documents and leverage the information to answer the questions, we introduce a novel prompting strategy for QA tasks with MixRAG called RECAP (Restate, Extract, Compute, Answer, Present), inspired by prompt strategies such as CoT (CoT) and EEDP (srivastava-etal-2024-evaluating) that involve step-by-step analytical reasoning, as well as PoT (PoT) which generates executable programs to solve complex mathematical operations. RECAP guides the LLM through a five-step linear process for reasoning and answering. Among these five steps, the Compute step plays a central role in improving answer quality, as it explicitly directs the model to reason and output intermediate logical steps or calculations when applicable.
@@
Formally, as defined in Section 3.1, the Multi-Step Reasoning Module is represented as $A=\text{Inference}(Q,D^{*})$, where $A$ is the final answer produced by the RECAP strategy. The five-step workflow of RECAP is as follows:
  1. (1)
    Restate the question: Clarify the question requirements and identify its type (e.g., numerical, fact-based, multi-hop).
  2. (2)
    Extract relevant data: Extract and list all content from $D^{*}$ that is directly relevant to the question.
  3. (3)
    Perform logical reasoning: Guide the LLM to perform explicit reasoning or computations based on the extracted information. If applicable, the LLM is prompted to write out the full formula used in the computation. This step is crucial for improving accuracy in questions requiring logical operations or quantitative analysis.
  4. (4)
    Generate final answer: Generate a final answer based on the above step. If needed, an external calculator uses the formula provided to validate or refine the answer. A rule-based mechanism then selects the more appropriate of the two.
  5. (5)
    Present selected answer: Return the selected answer in a clear and concise format.
Unlike methods like Self-Consistency (self-consistency), which improve accuracy by generating multiple reasoning paths, RECAP adopts a single-turn dialogue strategy, with its logical reasoning and formula provision originating from the same reasoning path, offering advantages in token efficiency for the LLM.
@@
Table 2. Comparison of RECAP in MixRAG Document-QA Performance (EM Score%) with Various Baselines
Models
Method GPT-4o GPT-4o mini Gemini-2.0 Qwen-Plus Qwen2.5-32B Qwen2.5-7B Mistral-Nemo Llama3.1-7B
Direct 35.34 23.55 26.64 34.83 21.80 8.02 10.03 15.79
CoT 46.87 36.84 38.01 50.20 48.37 31.58 25.06 25.31
EEDP 47.37 37.59 40.37 51.23 48.62 28.57 27.57 20.80
PoT 61.90 55.64 51.84 60.45 55.89 37.84 34.09 29.32
RECAP 64.66 58.15 60.25 64.45 64.16 45.36 44.11 40.60
5. Experiment
5.1. Experimental Setup
To evaluate the effectiveness of Heterogeneous Document RAG, we compare MixRAG against several representative baselines. For the retrieval task, we include several representative baselines:
  • LangChain (mavroudis2024langchain): A semi-structured retrieval pipeline that processes tabular and textual data by summarizing content with LLMs and conducting similarity-based retrieval.
  • SelfRAG (asai2023selfrag): A self-reflective retrieval framework that dynamically retrieves passages and uses reflection tokens to iteratively assess and refine outputs based on the model’s own feedback.
  • RAG-Critic (dong-etal-2025-criticrag): A framework based on a hierarchical error system that aligns a specialized error-critic model to guide an agentic workflow, facilitating error-driven self-correction.
  • Table Retrieval (chen-etal-2024-table): A method tailored for table retrieval that discovers join relationships and applies mixed-integer programming-based re-ranking to optimize both table-query and table-table relevance.
For the QA task, we consider several widely used reasoning baselines demonstrated strong performance in answering structured and semi-structured questions:
  • COT (CoT): A reasoning-based prompting strategy that guides the model to generate intermediate reasoning steps, improving its ability to solve complex problems.
  • EEDP (srivastava-etal-2024-evaluating): A multi-step reasoning framework that decomposes the answer process into explanation, refinement, and decision phases to enhance accuracy.
  • POT (PoT): A technique that prompts the LLM to produce a structured program for reasoning, which is then executed externally to yield precise results, especially for numerical problems.
We adopt HiT@K and Exact Match (EM) as evaluation metrics for retrieval performance and QA accuracy, respectively. To evaluate efficiency, we employ token usage and wall-clock time as our metrics. Further details are available in Appendix B. We employ GPT-4o throughout our pipeline to align with model used by baselines, including heterogeneous document representation, LLM-based retrieval, and RECAP execution. For embedding-based retrieval, we use OpenAI’s text-embedding-3-large model as the encoder. According to Appendix C, we set the parameters of MixRAG to $n=40$ and $m=60$ for ensemble retrieval. For the QA instance RECAP, the number of examples for all prompt strategies is uniformly set to 1. We evaluate the RECAP instance using GPT-4o, GPT-4o mini, Gemini-2.0-Flash and three open-source large language models, Qwen2.5-32B-Instruct, Mistral-Nemo-Instruct-2407, Llama-3.1-8B-Instruct.
5.2. Comparison with Baselines
In this section, we present a comprehensive analysis of MixRAG compared with several strong retrieval-augmented baselines, across the DocRAGLib dataset to evaluate effectiveness and efficiency.
5.2.1. Effectiveness Comparison
Retrieval Performance Comparison with Baselines. Table 1 presents the retrieval performance of MixRAG compared to several baselines across three different LLM backbones. Across all HiT@K metrics (K=1, 3, 5, 10), MixRAG consistently achieves the best results, reaching a HiT@1 score of 54.1% on GPT-4o, significantly outperforming the strongest baseline (Table Retrieval) by 46%. The fine-grained row-and-column-level summarization, combined with the two-stage retrieval strategy, achieves substantial improvements in retrieval accuracy for complex heterogeneous documents.
Document-QA Accuracy Comparison with Baselines. Using RECAP as the QA instance module for each retrieval method, we evaluate document-level QA accuracy based on EM scores. According to the EM results in Table 1, MixRAG also achieves the highest QA accuracy, with an EM score of 32.28% which corresponds to a 27.1% improvement over the second-best baseline, Table Retrieval.
In addition, we separately evaluate RECAP on the QA task to assess its standalone effectiveness compared to other table reasoning baselines on DocRAGLib. As shown in Table 2, RECAP consistently outperforms all competing methods across five LLMs, demonstrating its strong capability in handling heterogeneous document QA. Notably, by leveraging external calculation functions for arithmetic reasoning, RECAP improves the accuracy of computation-intensive questions while avoiding the instability often observed in program-executed strategies like PoT. Furthermore, its decomposition-based prompting strategy enables multi-step reasoning, especially in scenarios with heterogeneous document.
These results highlight the effectiveness of our retrieval and reasoning framework in improving answer accuracy for heterogeneous documents retrieval in downstream QA tasks.
5.2.2. Efficiency Comparison
We evaluate the efficiency of MixRAG in terms of token consumption and wall-clock time using GPT-4o. As detailed in Table 3, MixRAG strikes a balance between cost and performance. While it naturally incurs higher latency than single-step baselines (e.g., Standard RAG) due to its multi-hop reasoning mechanism, it is significantly more efficient than other iterative frameworks. Specifically, MixRAG reduces generation tokens by approximately 41% compared to Self-RAG and 59% compared to RAG-Critic. Notably, MixRAG achieves the lowest preprocessing token costs among all retrieval baselines, as well as lower generation token costs compared to iterative baselines, demonstrating its cost-efficiency. Overall, MixRAG achieves a superior trade-off between effectiveness and efficiency, delivering high reasoning performance while maintaining significantly low computational overhead.
Table 3. Comparison of Token Cost and Time on DocRAGLib
Method Pre. Tok. Gen. Tok. Time(s)
Standard RAG 0 4614.5 5.26
Table Retrieval 1339.6 5371.9 3.81
LangChain 5212.5 6944.4 3.37
Self-RAG 2551.3 24165.9 15.47
RAG-Critic 4917.3 34989.8 23.89
MixRAG 1108.3 14189.4 11.76
5.3. Ablation Study
Table 4. Performance Evaluation of MixRAG with Different Table Summary Level
Heterogeneous Document Representation HiT@K(%) EM*(%)
Table Summary Level K=1 K=3 K=5 K=10
Table Level Summary 36.27 48.05 59.63 74.28 23.83
General RCL Summary 39.84 55.02 68.03 79.81 26.25
H-RCL Summary 54.10 72.44 76.03 86.89 32.28
@@ 5.3. Ablation Study
Impact of Heterogeneous Document Representation Strategy in MixRAG. To evaluate the impact of different table summarization methods in heterogeneous document representation module, we compare three variants: Table Level Summary, General RCL Summary, and H-RCL Summary. As shown in Table 4, the H-RCL Table Summary strategy achieves the best performance among all heterogeneous document representation strategies. The General RCL Summary ranks second, showing clear gains over the traditional Table Level Summary by providing more explicit representations of rows and columns. H-RCL improves HiT@1 by 47% and EM by 28% compared to the Table Level Summary, demonstrating the importance of capturing multi-level table structures.
The Table Level Summary performs poorly because it tends to over-compress complex tabular data into flat textual descriptions. It sacrifices fine-grained details and row-column relationships, making it difficult for the retriever to locate precise evidence. The General RCL strategy ranks second. It provides a clearer representation of rows and columns, reducing ambiguity and improving retrieval accuracy, which in turn enhances QA performance by enabling more precise document matching. However, it treats tables as flat grids and struggles to capture the nested dependencies found in hierarchical tables. Building on this, the H-RCL strategy addresses these limitations by further capturing multi-level table structures and incorporating hierarchical path information. This design preserves the full lineage of cell values, significantly boosting both retrieval accuracy and the model’s ability to perform complex reasoning over structured data. Overall, these results demonstrate the effectiveness of the H-RCL design in improving both retrieval and reasoning capabilities for heterogeneous documents.
Table 5. Ablation Study Result of MixRAG
HiT@1(%) EM(%)
Ours (MixRAG) 54.10 32.38
@@
   only Embedding 47.34 27.77
   only BM25 35.45 21.00
@@
   only Embedding 27.87 21.31
   only BM25 15.57 13.11
- w/o R of RECAP 54.10 31.45
- w/o E of RECAP 54.10 30.63
- w/o C of RECAP 54.10 25.61
@@
Impact of Retrieval and QA Module. To investigate the effectiveness of each key component in MixRAG, we conduct the ablation study using DocRAGLib dataset. Table 5 shows the results of the Retrieval and QA module ablation study. Replacing Ensemble Retrieval with either Embedding or BM25 significantly reduces HiT@1, with Embedding achieving better results due to its stronger semantic alignment capabilities. Further removing LLM-based retrieval leads to an additional decline of 19% and 20% in HiT@1 for Embedding and BM25, respectively, confirming the advantage of the two-stage retrieval framework over single-stage methods. Moreover, disabling external calculation in RECAP results in a noticeable drop in EM, underscoring the importance of precise formula-based reasoning for improving answer accuracy. Overall, these results validate the effectiveness of each component within MixRAG in enhancing both retrieval precision and answer accuracy.
5.4. Scalability Analysis